OpenStudy (anonymous):

Tabitha is trying to find the equation of a line perpendicular to y = 1 over 2x − 5 in slope-intercept form that passes through the point (2, −7). Which of the following equations will she use?

OpenStudy (anonymous):

a) y − (−7) = 1 over 2(x − 2) b) y − 2 = 1 over 2(x − (−7)) c) y − (−7) = −2(x − 2) d) y − 2 = −2(x − (−7))

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you been at this one for a while huh?

OpenStudy (anonymous):

yes! can't seem to figure it out :(

OpenStudy (anonymous):

ok lets go slow (but not that slow) the slope of \[y=\color{red}{\frac{1}{2}}x-5\] is ?

OpenStudy (anonymous):

hint: it is red

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

right and the perpendicular line with have a slope that is the "negative reciprocal" i.e. flip it and change the sign

OpenStudy (anonymous):

2/1?

OpenStudy (anonymous):

-2/1*

OpenStudy (anonymous):

right, which, on planet earth, we usually just call \(-2\)

OpenStudy (anonymous):

haha right

OpenStudy (anonymous):

now we got the slope the point is \((2,-7)\) and they want you to plug all this in to the "point - slope" formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

where you see \(y_1\) put \(-7\) where you see \(m\) put \(-2\) and where you see \(x_1\) put \(2\)

OpenStudy (anonymous):

so would it be c?

OpenStudy (anonymous):

yes, it is always C how hard was that? ok?

OpenStudy (anonymous):

thank you so much!